OpenStudy (anonymous):

A light ray travels from water (n = 1.33) to crown glass (n = 1.52) and bends at a 30º angle to the normal. What is the incident angle for the light from the water? Answer with three significant digits. **I tried Snell's Law and no matter what I do, I didn't get the correct answer, help me :(

OpenStudy (theeric):

It seems like you have all the information for Snell's Law! \(n_w\sin\theta_w=n_g\sin\theta_g\) \(\implies \theta_w=\arcsin\left(\dfrac{n_g\sin\theta_g}{n_w}\right)\)

OpenStudy (theeric):

 That doesn't seem too bad :) What are you getting?

OpenStudy (theeric):

Good luck! If you're using a calculator, make sure you have it set to degrees!